# Artworks Reimagined: Exploring Human-AI Co-Creation through Body Prompting

### v1

@@ Abstract.
Image generation using generative artificial intelligence is a popular activity. However, it is almost exclusively performed in the privacy of an individual’s home via typing on a keyboard. In this article, we explore body prompting as input for image generation. Body prompting extends interaction with generative AI beyond textual inputs to reconnect the creative act of image generation with the physical act of creating artworks. We implement this concept in an interactive art installation, Artworks Reimagined, designed to transform artworks via body prompting. We deployed the installation at an event with hundreds of visitors in a public and private setting. Our results from a sample of visitors ($N=79$) show that body prompting was well-received and provides an engaging and fun experience. We identify three distinct patterns of embodied interaction with the generative AI and present insights into participants’ experience of body prompting and AI co-creation. We provide valuable recommendations for practitioners seeking to design interactive generative AI experiences in museums, galleries, and other public cultural spaces.
human-AI interaction, embodied interaction, image generation, generative AI, co-creation
copyright: rightsretainedjournalyear: 2024ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Gestural input
Figure 1. The study setup of the interactive art installation, Artworks Reimagined, guides participants through four application screens (top) in which participants (1) give consent, (2) select an artwork, (3) body prompt via posing in front of a camera, and (4) exit the camera booth. This is followed by an interview in which the results are revealed (5).
[…]
The study setup and user flow through four application screens (top) in which users give consent (1), select an artwork (2), pose (3), and exit the camera booth via an end screen (4). At the event, this was followed by an interview in which the results were revealed (5).
@@ 1. Introduction
Image generation with artificial intelligence (AI) has gained popularity, with millions of practitioners exploring text-to-image generation systems for recreation and creative exploration (Oppenlaender, 2022; McCormack et al., 2024; Kahil, 2023; Sanchez, 2023). However, text-to-image generation is almost exclusively conducted in the privacy of an individual’s home, and limited to typing on a computer keyboard. We argue the current focus of image generation on typing and discrete prompts by one solitary creator limits opportunities for engaging experiences, due to four key constraints of text-to-image generation: literacy, input control, social creativity, and the setting.
Limited by literacy: Frequently regarded as an instrumentalist use of technology, text-to-image generation is a process in which AI is used as a tool to visually translate concepts or ideas (Coeckelbergh, 2023). For this translation to be effective, English literacy is required. Further, some image generators require the use of specific keywords to produce high-quality outputs (Oppenlaender, 2023), which demands expertise in “prompt engineering” (Oppenlaender, 2024, 2022, 2023; Oppenlaender et al., 2023a) and familiarity with art concepts, such as styles and media (Oppenlaender, 2023). These language-based requirements exclude certain populations, such as the illiterate, young children, and all those who are not familiar with the English language.
Limited input control and expressivity: The level of control over text-to-image generation is limited by language (McCormack et al., 2023). Prompt engineering is tedious and it is not possible to fully control text-to-image generation with discrete prompts (Oppenlaender et al., 2023a), leaving much of the initial image generation to randomness (Hauhio et al., 2023). In practice, the first generated image is often one step in a longer creative process that involves many tools, such as image editors, image-to-image generation, inpainting, and others (Oppenlaender, 2024; Palani and Ramos, 2024; Sanchez, 2023). ControlNet was a remarkable breakthrough technology (Zhang et al., 2023), allowing fine-grained control over image generation via human poses and other input forms. A pose is a skeleton-like structure with key points that correspond to major joints and body parts. ControlNet gained widespread popularity among AI art practitioners. However, rarely do these poses originate from the practitioners’ own bodies. Instead, poses are transferred from existing images, taken from openpose datasets, or created with tedious manual editing. In addition to questions about copyright (Samuelson, 2023; Oppenlaender et al., 2023b; Samuelson, 2024), this “raises issues of embodiment, materiality, and material agency that traditionally come into play when making a visual artwork” (McCormack et al., 2023).
Limited collaboration and social creativity: Contemporary literature views creativity as a social activity (Frich et al., 2019; Amabile, 1983), with the concept of the lone creator being dispelled as a myth (Lemley, 2012). Yet, image production with generative AI is almost exclusively completed by a solitary creator. The social and performative aspects of art creation are lost due to the privacy of the image generation setting and the constrained text input modality. Collaborative websites, such as Artbreeder [^1] https://www.artbreeder.com, have demonstrated that collaborative image generation can be an engaging experience, with wide audience appeal.
Limited suitability for public settings: Confined to the personal and narrow triangular interaction space between user, keyboard, and computer display, the practice of text-to-image generation does not lend itself well to public settings, such as Galleries, Libraries, Archives, and Museums (GLAM). For such public settings, input modalities should be dynamic and intuitive, offering glimpses of the generative process to the audience and building anticipation for the final image. Typing on a keyboard in front of a computer screen is not expressive, in this regard, and does not provide an engaging audience experience.
Addressing these limitations could make image generation more inclusive and accessible to a wider audience (Kinnula et al., 2021). Prior work aimed to break free from these four limitations by exploring audience participation in image generation in a museum context. Kun et al.’s interactive GenFrame allowed museum visitors to shape the image generation process via three knobs controlling different parameters of the image generation process (Kun et al., 2024a, b). However, this interaction and the input modality are limited, leaving much of the agency in the act of co-creation to the AI system, not the user.
In this paper, we explore body prompting as a novel input modality for image generation. A body prompt is an embodied input for the image generation system in form of a human pose. Body prompting extends image generation interaction beyond textual inputs to reconnect the creative act of image generation with the expressive physical act of art creation. We enable an audience in a public setting to “reimagine” an existing artwork with their body pose, thereby creating their own version of it. To this end, we designed and implemented Artworks Reimagined, an interactive art installation conceived for the European Researchers’ Night 2023 that, using generative AI, transforms and reinterprets classical artworks via body prompting.
@@
We evaluate body prompting in a user study conducted in a real-world setting, specifically, at a public event with hundreds of visitors. However, some people may be shy and feel inhibited when performing in public, and society is biased against introverts (Cain, 2012). For this reason, we designed and tested two separate body prompting settings: a public staging area where people could body-prompt in front of an audience, and a private booth where people could body-prompt and view the results hidden from the public eye (cf. Figure 2). Our investigation focused on observing participants’ interactions, behavior, and body prompting strategies, followed by a semi-structured interview. This study is exploratory, emphasizing how people naturally engage with generative technology via body prompting when encountering it in a public setting without specific instructions. Our study is guided by the following research questions:
@@
    Generated images: What images do participants create with body prompting? Is there a difference in images between the public and private setting?
@@
    User experience: How do participants experience body prompting? And how do they experience AI co-creation via body prompting?
@@
    Participant behavior: What are the decisions and preferences of participants in regard to body prompting? What goals do participants pursue when body prompting?
@@
    Body prompting in public: How do personality traits, as measured by the “Big Five” (Digman, 1990), affect body prompting and the use of the installation?
After reviewing related work in Section 2, we present the design of Artworks Reimagined and the study methodology in Section 3. We then provide an overview of the images generated and a synthesis of observations from the event in Section 4. We discuss our findings and provide design recommendations for practitioners and researchers in Section 5. We conclude in Section 6.
@@ 2. Related Work
Image generation is predominantly performed via typing prompts on a keyboard in a private space. Therefore, in the vast majority of cases, prompting is “an isolated and individual experience” (Mahdavi Goloujeh et al., 2024). In this section, we focus on related work that sought to break free from this reigning interaction paradigm to develop new inputs and bring the activity of image generation into the public sphere. We first review research on interaction with public displays (Section 2.1), and then examine embodied experiences (Section 2.2) and generative AI in public installations (Section 2.4).
@@ 2.1. Interaction with Public Displays
Public displays offer opportunities to study how users interact with technology. However, public displays bring unique challenges, such as display blindness (Müller et al., 2009), interaction blindness (Hosio et al., 2016), and the first-click problem (Kukka et al., 2013). The latter arises when users are unsure how to initiate interaction with a public display. In this section, we focus on challenges related to the public setting of our study.
Privacy concerns require consideration. Brudy et al. identified ‘shoulder surfing’ – where unauthorized individuals may view sensitive information – as a pertinent risk associated with public displays (Brudy et al., 2014). Further exploring privacy issues, Memarovic examined how the public’s concerns about photography in the vicinity of public displays influence user interaction (Memarovic, 2015). Their study emphasized the importance of transparent policies regarding the storage and usage of photos taken in these contexts, highlighting a general need for clear communication about the handling of users’ data. In a comparative study, Collier et al. addressed differences between private and public self-service technologies (Collier et al., 2014). Their findings suggest that perceptions of control and convenience significantly vary depending on the setting, with public technologies often viewed as less controllable and convenient than their private counterparts (Collier et al., 2014). This perception could influence user engagement and satisfaction with the public display in our study. Hosio et al. studied privacy in interactive public displays (Hosio et al., 2019; Alorwu et al., 2022). They found that people are surprisingly willing to volunteer highly sensitive personal information in public settings.
These studies reveal a complex landscape of user behavior and expectations when interacting with displays in public spaces. Addressing these concerns is crucial for designing an interactive public installation.
@@ 2.2. Interactive Experiences via Embodied Interaction
Embodied interaction can provide an engaging experience. This is particularly relevant in the context public cultural spaces, such as GLAM institutions. In museums, the contact with more traditional art forms, namely painting or sculpture, tends to occur passively through a cognitive engagement with the artworks. Digitally enriched museum experiences (Calvi and Vermeeren, 2023) open new opportunities to experience and interact with such artworks through embodied cognitive engagement via human-technology interaction. Gestures and posing are an engaging way of meaning making in the GLAM context (Steier et al., 2015; Steier, 2014; Fu et al., 2024). In the remainder of this section, we review prior work on embodied interaction in the context of GLAM institutions and other public settings.
Trajkova et al. explored engaging museum visitors with embodied interaction (Trajkova et al., 2020). Their study was aimed at human-data interaction, enabling people to explore data sets and interactive visualizations with gestures and body movements.
Oppenlaender and Hosio presented a system for museum visitors to provide feedback for artworks, including selfie photos and selfie videos (Oppenlaender and Hosio, 2019). The latter was perceived as a surprising way of interacting with an artwork and its artist. Kozinets et al. also explored museum selfies (Kozinets et al., 2017), finding selfies to be a way to construct narratives about oneself. However, their study did not feature embodied interaction with dynamically generated artifacts in the museum context.
Positive effects of interactive technological engagement with artworks have been reported. For example, aesthetic appreciation (including interest, intensity, pleasure, and learning) was higher in artworks with interactive elements compared to non-interactive physical objects (Jonauskaite et al., 2024). In Coeckelbergh’s framework, body prompting can be conceptualized as a performance-based approach to interaction and AI image generation (Coeckelbergh, 2023). Body prompting is a poietic performance within human-technology co-creation. The performance-based approach emphasizes the process of interaction in which novel artistic (quasi-)subjects and objects emerge and are produced in and by the processes instead of instruments and tools (Coeckelbergh, 2023).
@@ 2.3. Re-creating and Reinterpreting Artworks
Art museums are increasingly digitizing their collections and making them accessible on the internet thus fostering different reinterpretations and re-creations of the paintings (Barranha, 2018). Several experiments have shown that people enjoy re-creating and reinterpreting famous artworks. Launched at the height of the COVID-19 pandemic, The Getty Museum Challenge was a widely successful initiative, which invited the public to re-create an artwork with household items (Romero, 2020). In 2023, the Mauritshuis Museum in The Hague launched another challenge, which invited the public to reinterpret Vermeer’s “The Girl with a Pearl Earring” using different processes, including generative AI (Mauritshuis Museum, 2023). The AI-generated images were later exhibited in the Museum in place of Vermeer’s famous painting, in a digital loop display which featured a wide range of experiments produced with different media. In parallel to such institutional initiatives, generative AI has also been used by researchers to reinterpret artists’ self-portraits (Barranha, 2023).
These examples of re-creating and reinterpreting artworks suggest that new activities designed to explore art collections with the use of generative AI can foster participatory creative experiences. However, the Getty and Mauritshuis Museum challenges were completed from the participants’ homes, re-creating or reinterpreting the artwork either with body poses and household items or generative AI. In our work, we explored how visitors could be engaged to re-create or reinterpret an artwork with generative AI and their bodies at a public event.
@@ 2.4. Generative AI in Public Exhibitions
Generative AI has been explored in the context of public exhibitions. In this section, we review three interactive installations that used generative AI.
@@
GenFrame by Kun et al. (Kun et al., 2024a, b) is a more recent generative art piece designed for the museum context. GenFrame’s picture display gives the appearance of a traditional artwork, however the audience can control the image generation with three rotating knobs. Like our work, GenFrame was implemented with ControlNet and StableDiffusion. Nevertheless, the interaction affordances of GenFrame are limited by design. Our work explores body prompting as an expressive way of providing input to the image generation process.
3. Method
@@
This section first provides background information on the public event that was the context for our study and our motivation and design goals for Artworks Reimagined. We then describe the study design and procedure, research materials, data collection and analysis, and technical implementation.
3.1. Background and Design Goals

### v2

@@ Abstract.
Image generation using generative artificial intelligence has become a popular activity. However, text-to-image generation—where images are produced from typed prompts—can be less engaging in public settings since the act of typing tends to limit interactive audience participation, thereby reducing its suitability for designing dynamic public installations. In this article, we explore body prompting as input modality for image generation in the context of installations at public event settings. Body prompting extends interaction with generative AI beyond textual inputs to reconnect the creative act of image generation with the physical act of creating artworks. We implement this concept in an interactive art installation, Artworks Reimagined, designed to transform existing artworks via body prompting. We deployed the installation at an event with hundreds of visitors in a public and private setting. Our semi-structured interviews with a sample of visitors ($N=79$) show that body prompting was well-received and provides an engaging and fun experience to the installation’s visitors. We present insights into participants’ experience of body prompting and AI co-creation and identify three distinct strategies of embodied interaction focused on re-creating, reimagining, or casual interaction. We provide valuable recommendations for practitioners seeking to design interactive generative AI experiences in museums, galleries, and public event spaces.
generative AI, human-AI interaction, embodied interaction, image generation, co-creation, public displays, art installation
copyright: nonecopyright: rightsretainedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Gestural input
@@ 1. Introduction
Recent advances in generative artificial intelligence (AI) have fueled widespread interest in text-to-image generation systems, which millions of users now employ generative AI for creative exploration and recreation (Oppenlaender, 2022; Kahil, 2023; Oppenlaender, 2024; Sanchez, 2023). Yet these systems are typically used in private settings where users type discrete prompts on keyboards, a design that limits engagement in public contexts. This reliance on text-based input brings four key challenges—input control, literacy, social creativity, and physical setting constraints (detailed in Section 3.2)—that present barriers to accessibility and limit the potential for interactive, public experiences (Kinnula et al., 2021).
Building on earlier efforts to involve audiences in image generation in public settings, such as GenFrame (Kun et al., 2024a) in a museum setting, researchers have begun exploring alternative input methods for image generation. GenFrame, for instance, allows visitors to influence the image creation process through three rotating knobs, offering a novel interaction modality to image generation (Kun et al., 2024a, b). However, this approach restricts human agency because much of the creative decision-making remains with the AI system rather than with the user. Our work aims to address these limitations by exploring an interaction design that more effectively balances user input with algorithmic creativity, ultimately making image generation more inclusive and engaging in public events.
In this paper, we introduce body prompting as a novel input modality for image generation. Instead of relying solely on text, body prompting uses a person’s pose to guide the generative process, reconnecting image creation with the expressive physicality of traditional art-making. Our work, thus, touches on considerations of embodiment, materiality, and material agency that have long informed the creation of traditional visual art (McCormack et al., 2023).
Our goal is to empower audiences in public settings to “reimagine” existing artworks through their body movements, creating personalized reinterpretations. To support this vision, we designed and implemented an interactive art installation—Artworks Reimagined—which leverages generative AI to transform classical artworks based on users’ body prompts. The installation was deployed at the European Researchers’ Night 2023, a public event that attracts thousands of visitors.
@@
We evaluated body prompting in a field study at European Researchers’ Night, engaging hundreds of visitors. Acknowledging that some individuals may feel inhibited when performing in front of others (Cain, 2012), we implemented two interaction settings: a public staging area for body prompting in direct view of an audience, and a private booth where visitors could interact and view results without an on-looking audience (cf. Figure 3). Our investigation combines observations of participants’ interactions and strategies with follow-up semi-structured interviews ($N=79$), offering insights into how visitors naturally engage with generative technology via body prompting in a public context without structured guidance. Our evaluation of the system is guided by the following research questions:
@@
    Generated images: What images do visitors create with body prompting? Is there a difference in images between the public and private setting?
@@
    User experience: How do visitors experience body prompting? And how do they experience AI co-creation via body prompting?
@@
    Participant behavior: What are the decisions and preferences of visitors in regard to body prompting? What goals do visitors pursue when body prompting?
@@
    Body prompting in public: How do personality traits, as measured by the “Big Five” (Digman, 1990), affect body prompting and the use of the art installation in the private and public setting?
Our exploration of body prompting as an interaction method for generative AI art installations at public events indicates that this approach effectively engages both active participants and observers. Body prompting was perceived as an engaging and fun experience. Participants employed three distinct strategies: imitating the source artwork, reimagining it, and expressing themselves naturally. Notably, personality traits did not significantly influence the choice of posing strategy or setting, as both extroverts and introverts appreciated the experience.
We contribute to the field of Engineering Interactive Computing Systems (EICS) by engineering a novel interactive system that uses body prompting to drive generative AI. Our work describes a modular pipeline and software architecture based on T2I-Adapter with Stable Diffusion—combining ControlNet-based pose detection, style transfer, and CLIP-Interrogator for automated prompt generation—with dual web applications and cloud services (AWS Gateway, Lambda, S3) to generate personalized recreations and reimaginations of an existing digital artwork. This work advances engineering methods by providing a detailed architecture, workflow processes, and evaluation of user strategies that can be adapted to design and deploy co-creative interactive systems with generative AI in public settings. In addition, we make design recommendations for developing and refining interactive computing systems in dynamic, real-world settings.
The remainder of this work is structured as follows. After reviewing related work in Section 2, we present the design rationale motivating our design of Artworks Reimagined in Section 3. We present the installation’s overall concept, design, and implementation in Section 4. We then describe our evaluation of the system in Section 5. this includes an overview of the images generated, detailed analysis of interview responses, and a synthesis of observations from the event. We discuss our findings and provide design recommendations for practitioners and researchers in Section 6 and conclude in Section 7.
@@ 2. Related Work
Image generation is typically confined to private spaces where users type prompts, consisting of discrete words, on a keyboard. This results in an isolated, individual experience (Mahdavi Goloujeh et al., 2024). Recent research, however, has explored alternative input modalities and interaction paradigms to integrate image generation into public and social contexts. In this section, we review work that challenges the traditional text-based approach to image generation, and we review related work that presents challenges to its implementation. We first discuss challenges around interactive public displays (Section 2.1), then examine embodied experiences (Section 2.2), followed by a review of studies on re-creating and reinterpreting artworks (Section 2.3). Finally, we review selected studies that leveraged generative AI in public installations (Section 2.4).
@@ 2.1. Interaction with Public Displays
Public displays offer opportunities to study how users interact with technology. However, public displays bring unique challenges, such as display blindness (Müller et al., 2009), interaction blindness (Hosio et al., 2016), and the first-click problem (Kukka et al., 2013). These issues arise when users are unsure how to initiate interaction with a public display, or when they ignore the interactivity of the display entirely. In this section, we focus on challenges specifically related to the public setting of our study.
Privacy concerns require consideration in the design of public display installations. Brudy et al. (2014) identified ‘shoulder surfing’—where unauthorized individuals may view sensitive information—as a pertinent risk associated with public displays. Further exploring privacy issues, Memarovic (2015) examined how the public’s concerns about photography in the vicinity of public displays influence user interaction. Their study emphasized the importance of transparent policies regarding the storage and usage of photos taken in these contexts, highlighting a general need for clear communication about the handling of users’ data. In a comparative study, Collier et al. (2014) addressed differences between private and public self-service technologies. Their findings suggest that perceptions of control and convenience significantly vary depending on the setting, with public technologies often viewed as less controllable and convenient than their private counterparts (Collier et al., 2014). This perception could influence user engagement and satisfaction with the public display in our study. Hosio et al. studied privacy in interactive public displays (Hosio et al., 2019; Alorwu et al., 2022). They found that people are surprisingly willing to volunteer highly sensitive personal information in public settings.
These studies reveal the complexity of user behavior and expectations when interacting with displays in public spaces. Addressing these concerns is crucial for designing an interactive public installation.
@@ 2.2. Interactive Experiences via Embodied Interaction
Embodied interaction can provide an engaging experience. This is particularly relevant in the context of public cultural spaces, such as GLAM institutions (Galleries, Libraries, Archives, and Museums). In museums, the contact with more traditional art forms, namely painting or sculpture, tends to occur passively through a cognitive engagement with the artworks. Digitally enriched museum experiences (Calvi and Vermeeren, 2023) open new opportunities to experience and interact with such artworks through embodied cognitive engagement via human-technology interaction. Gestures and posing are an engaging approach to meaning-making in the GLAM context (Steier et al., 2015; Steier, 2014; Fu et al., 2024). In the remainder of this section, we review prior work on embodied interaction in the context of GLAM institutions and other public settings.
Trajkova et al. (2020) explored engaging museum visitors with embodied interaction. Their study was aimed at human-data interaction, enabling visitors to explore data sets and interactive visualizations with gestures and body movements. Oppenlaender and Hosio (2019) presented a system for museum visitors to provide interactive feedback for artworks, including taking selfie photos and selfie videos. The latter was perceived as a surprising way of interacting with an artwork and its artist. Kozinets et al. (2017) also explored museum selfies, finding selfies to be a way to construct narratives about oneself. However, their study did not feature embodied interaction with dynamically generated artifacts in the museum context.
Positive effects of interactive technological engagement with artworks have been reported. For example, aesthetic appreciation (including interest, intensity, pleasure, and learning) was higher in artworks with interactive elements compared to non-interactive physical objects (Jonauskaite et al., 2024). In Coeckelbergh’s framework (Coeckelbergh, 2023), body prompting can be conceptualized as a performance-based approach to interaction and AI image generation. Body prompting is a poietic performance within human-technology co-creation. The performance-based approach emphasizes the process of interaction in which novel artistic (quasi-)subjects and objects emerge and are produced in and by the processes instead of instruments and tools (Coeckelbergh, 2023).
@@ 2.3. Re-creating and Reinterpreting Artworks
Art museums are increasingly digitizing their collections and making them accessible on the internet thus fostering different reinterpretations and re-creations of the paintings (Barranha, 2018). Several experiments have shown that people enjoy re-creating and reinterpreting famous artworks. Launched at the height of the COVID-19 pandemic, The Getty Museum Challenge (Romero, 2020) was a widely successful initiative, which invited the public to re-create an artwork with household items. In 2023, the Mauritshuis Museum in The Hague launched another challenge, which invited the public to reinterpret Vermeer’s famous painting “The Girl with a Pearl Earring” using different processes, including generative AI (Mauritshuis Museum, 2023). The AI-generated images were later exhibited in the Museum in place of Vermeer’s famous painting, in a digital loop display which featured a wide range of experiments produced with different media. In parallel to such institutional initiatives, generative AI has also been used by researchers to reinterpret artists’ self-portraits (Barranha, 2023).
These examples of re-creating and reinterpreting artworks suggest that new activities designed to explore art collections with the use of generative AI can foster participatory creative experiences. However, the Getty and Mauritshuis Museum challenges were completed from the participants’ homes, re-creating or reinterpreting the artwork either with body poses and household items or generative AI. In our work, we explore how visitors can be engaged to re-create or reinterpret an artwork with their bodies and generative AI at a public event.
@@ 2.4. Generative AI in Public Exhibitions
Generative AI has been explored in the context of exhibitions, GLAM settings, and other public spaces. In this section, we review three selected interactive installations that used generative AI for engaging audiences.
@@
GenFrame by Kun et al. is a recent related study featuring a generative art piece designed for the museum context (Kun et al., 2024a, b). GenFrame’s framed display gives the appearance of a traditional artwork, however the audience can control the image generation with three rotating knobs. Like our work, GenFrame was implemented with ControlNet and StableDiffusion. Nevertheless, the interaction affordances of GenFrame are limited by design. Our work explores body prompting as an expressive way of providing input to the image generation process.
3. System Design
@@
In this section, we first provide background information on the public event that was the context for our study and our motivation and design goals for Artworks Reimagined in Section 3.1. We then describe four key challenges based on our formative study in Section 3.2 and the design goals of our system in Section 3.3.
3.1. Formative Study